### v2

# Log-Lipschitz and Hölder regularity imply smoothness for complex analytic sets

Abstract.
In this paper, we prove metric analogues, in any dimension and in any co-dimension, of the famous Theorem of Mumford on smoothness of normal surfaces and the beautiful Theorem of Ramanujam that gives a topological characterization of $\mathbb{C}^{2}$ as an algebraic surface. For instance, we prove that a complex analytic set that is log-Lipschitz regular at 0 (i.e., a complex analytic set that has a neighbourhood of the origin which bi-log-Lipschitz homeomorphic to an Euclidean ball) must be smooth at 0. We prove even more, we prove that if a complex analytic set $X$ such that, for each $10<\alpha<$ 1, $(X,0)$ and $(\mathbb{R}^{k},0)$ are bi-$\alpha$-Hölder homeomorphic, then $X$ must be smooth at 0. These results generalize the Lipschitz Regularity Theorem, which says that a Lipschitz regular complex analytic set must be smooth. Global versions of these results are also presented here and, in particular, we obtain a characterization of an affine linear subspace as a pure-dimensional entire complex analytic set.
Key words and phrases:
Log-Lipschitz regularity, Lipschitz regularity, Hölder regularity, Smoothness of analytic sets, Theorem of Mumford
2010 Mathematics Subject Classification:
14B05; 32S50
The author was partially supported by CNPq-Brazil grant 310438/2021-7. This work was supported by the Serrapilheira Institute (grant number Serra – R-2110-39576).
1. Introduction
In 1961, D. Mumford in his famous article [25] proved that a complex algebraic surface, with normal singularity at $x_{0}$, is smooth at $x_{0}$ if and only if its link at $x_{0}$ is simply connected. Since in the case of a complex normal surface $X$, its link at $x_{0}\in X$ is simply connected if and only if $X$ is topologically regular at $x_{0}$ (i.e., there is an open neighbourhood $U$ of $x_{0}$ such that $X\cap U$ is homeomorphic to some open euclidean ball), we can state the Theorem of Mumford as the following:
Theorem 1.1 (Theorem of Mumford).
Let $X\subset\mathbb{C}^{n}$ be a complex algebraic set with a normal singularity at $x_{0}$. Then $X$ is smooth at $x_{0}$ if and only if $X$ is topologically regular at $x_{0}$.
Unfortunately, this result does not hold true in dimensions greater than 2, as it was proved by F. Brieskorn in [5] (see also [6]). For instance, he proved that the hypersurface $X=\{(x,y,z,w)\in\mathbb{C}^{4};x^{2}+y^{2}+z^{2}=w^{3}\}$, which has a normal singularity at 00, is topologically regular at 00, but is not smooth at 00. Thus, in order to obtain a result like the Theorem of Mumford in higher dimension, we have to impose more conditions. As another reference of works in the same line, we also have D. Prill [26] where it was proved that any d-dimensional complex algebraic cone with trivial i-th homotopy group of the link, for all $20\leq i\leq 2d-$ 2, must be a linear affine subspace. One more result of the type of Theorem of Mumford was obtained separately by N. A’Campo in [1] and Lê D. T. in [21], which can be stated as follows: Let $(X,x_{0})$ be a codimension one germ of complex analytic subset of $\mathbb{C}^{n}$. If there is a homeomorphism $\varphi\colon(\mathbb{C}^{n},X,x_{0})\to(\mathbb{C}^{n},\mathbb{C}^{n-1}\times\{0\},x_{0})$, then $X$ is smooth at $x_{0}$.
However, the Theorem of A’campo-Lê does not hold, in general, for complex analytic sets that are topological submanifolds in codimension greater than 1. For instance, the cusp $C=\{(x,y,z)\in\mathbb{C}^{3};x^{3}=y^{2}$ and $z=0\}$ is a topological submanifold of $\mathbb{C}^{3}$, but is not smooth at 00.
Recently, it was presented two different approaches for a type of Theorem of Mumford in higher dimension: one from the contact geometry point of view and the other one from the Lipschitz geometry point of view.
From the contact geometry point of view, M. McLean in [23] showed that: if a complex analytic set $A$ has a normal isolated singularity at 00 and its link at 00 (with its canonical contact structure) is contactomorphic to the link of $\mathbb{C}^{3}$ (the standard contact sphere), then $A$ is smooth at 00.
From the Lipschitz geometry point of view, the author in his thesis [28] (see also [29] and [4]) proved the following:
Theorem 1.2 (Lipschitz Regularity Theorem).
A complex analytic set $X$ is smooth at $x_{0}$ if and only if $X$ is, around $x_{0}$, bi-Lipschitz homeomorphic to an open Euclidean ball.
Around a decade ago the author was presented, in a particular conversation, to the following problem:
Problem 1.
Let $X\subset\mathbb{C}^{n}$ be a complex analytic set of dimension $d$. If $X$ is log-Lipschitz regular at 00, then is $X$ smooth at 00?
In order to know, $h$ is log-Lipschitz if there exists a positive constant $C$ such that $\|h(x)-h(y)\|\leq C\|x-y\|\big{|}\log\|x-y\|\big{|}$ for all $x,y\in X$ such that $20<\|x-y\|<1/$ 2. Thus, we say that $X$ is log-Lipschitz regular at 00 if there are an open neighbourhood $U$ of 00 and a homeomorphism $h\colon X\cap U\to B_{r}^{k}(0)$ such that $h(0)=0$ and $h$ and $h^{-1}$ are log-Lipschitz. Similarly, we define the notions of $\alpha$-Hölder and Lipschitz regularities at 00 (see the definitions of the different regularities in Subsection 2.1).
In this article, we answer affirmatively Problem 1. Indeed, we obtain the following result of smoothness of analytic sets, which gives even more than a positive answer to Problem 1.
Theorem 3.1.
Let $X\subset\mathbb{C}^{n}$ be a complex analytic set. Then the following statements are equivalent:
  • (1)
    $X$ is $\alpha$-Hölder regular at 00 for all $\alpha\in(0,1)$;
  • (2)
    $X$ is strongly Hölder regular at 00;
  • (3)
    $X$ is log-Lipschitz regular at 00;
  • (4)
    $X$ is Lipschitz regular at 00;
  • (5)
    $X$ is smooth at 00.
Note that in the very particular case when $\dim X=1$, Theorem 3.1 is a consequence of the main result in [14] (see some related studies in [10] and [15]).
Theorem 3.1 is sharp in the following sense: for any $\alpha\in(0,1)$ and any positive integer $d$, there is a complex algebraic hypersurface $X\subset\mathbb{C}^{d+1}$ that is $\alpha$-Hölder regular at 00, but is not smooth at 00 (see Proposition 3.2).
The following example shows that one cannot obtain a result like Theorem 3.1 for sets in general.
Example 1.3.
Let $f\colon\mathbb{R}\to\mathbb{R}$ be the function given by
$f(x)=\left\{\begin{array}[]{ll}|x||\log|x||,&\mbox{ if }x\not=0,\\ 0,&\mbox{ if }x=0.\end{array}\right.$
Then $X=Graph(f)$ is log-Lipschitz regular at 00, but is not Lipschitz regular at 00. Indeed, the restriction of the projection on the first coordinate $\pi\colon X\to\mathbb{R}$ is a bi-log-Lipschitz homeomorphism.
It is clear that Theorem 3.1 is a generalization of the Lipschitz Regularity Theorem. However, Example 1.3 also shows that the proof of the Lipschitz Regularity Theorem cannot be used to prove Theorem 3.1. Indeed, the main ingredients in the proof of the Lipschitz Regularity Theorem were the following:
  • (1)
    The bi-Lipschitz invariance of the tangent cones in [28, Teorema 4.1.1] (see also [29, Theorem 3.2], [17, Theorem 3.6] and [34, Theorem 3.1]), which says that if two germs of definable sets in an o-minimal structure on $\mathbb{R}$ are bi-Lipschitz homeomorphic, then their tangent cones are bi-Lipschitz homeomorphic too (see the definition of tangent cone in Definition 2.4);
  • (2)
    Lipschitz regularity implies Lipschitz normally embeddedness, i.e., a set that is Lipschitz regular at 00 must be Lipschitz normally embedded (or quasi-convex) at 00 (see Definition 2.3).
However, there are germs of sets $(A,a)$ and $(B,b)$ that are bi-log-Lipschitz homeomorphic, but their tangent cones are not homeomorphic. For instance, in Example 1.3, $C(X,0)=\{(0,y)\in\mathbb{R}^{2};y\geq 0\}$ and $C(\mathbb{R},p)=\mathbb{R}$, for any $p\in\mathbb{R}$. Moreover, there are sets that are log-Lipschitz regular at 00, but are not Lipschitz normally embedded at 00. For instance, the set $X$ in Example 1.3 is log-Lipschitz regular at 00, but is not Lipschitz normally embedded at 00.
In order to know more about the o-minimal geometry, see, for instance, [7] and [9].
Still in this article, we also present a global version of Theorem 3.1, which can be seen as a Lipschitz analogous of the following beautiful topological characterization of $\mathbb{C}^{2}$ as an algebraic surface, presented by Ramanujam in [27, Theorem].
Theorem 1.4 (Theorem of Ramanujam).
Let $X$ be an affine smooth complex algebraic surface that is contractible and simply connected at infinity. Then $X$ is isomorphic to $\mathbb{C}^{2}$ as an algebraic variety.
Nowadays, the Theorem of Ramanujam is equivalent to say that any affine smooth complex algebraic surface $X\subset\mathbb{C}^{n}$, which is contractible and $X\setminus B_{r}^{2n}(0)$ is homeomorphic to $\mathbb{C}^{2}\setminus B_{r}^{4}(0)$ for all big enough $r>0$, must be isomorphic to $\mathbb{C}^{2}$ as an algebraic variety.
Another result in this line was presented by Ahern and Rudin in [3]. To be more precise, Ahern and Rudin in [3] defined the notion of a set to be $C^{1}$-smooth at infinity.
Definition 1.5.
A set $V\subset\mathbb{R}^{n}$ is $C^{1}$-smooth at infinity if $\iota(V\setminus\{0\})\cup\{0\}$ is a $C^{1}$ submanifold around 00, where $\iota\colon\mathbb{R}^{n}\setminus\{0\}\to\mathbb{R}^{n}\setminus\{0\}$ is the mapping given by $\iota(x)=\frac{x}{\|x\|^{2}}$.
With this notion, Ahern and Rudin in [3] proved the following very interesting result:
Theorem 1.6 (Thereom II in [3]).
Let $V\subset\mathbb{C}^{n}$ be a complex analytic set. Then $V$ is $C^{1}$ smooth at infinity if and only if $V$ is the union of an affine linear subspace of $\mathbb{C}^{n}$ and a (possibly empty) finite set.
Similarly, we define also the following:
Definition 1.7.
A set $V\subset\mathbb{R}^{n}$ is:
  • $\alpha$-Hölder smooth at infinity if $\iota(V\setminus\{0\})\cup\{0\}$ is $\alpha$-Hölder regular at 00;
  • strongly Hölder smooth at infinity if $\iota(V\setminus\{0\})\cup\{0\}$ is strongly Hölder regular at 00;
  • log-Lipschitz smooth at infinity if $\iota(V\setminus\{0\})\cup\{0\}$ is log-Lipschitz regular at 00;
  • Lipschitz smooth at infinity if $\iota(V\setminus\{0\})\cup\{0\}$ is Lipschitz regular at 00.
Thus, for algebraic sets, the next result generalizes [3, Thereom II] and [11, Theorem 3.8] (see also [32, Theorem 1.6], [13, Corollary 4.3], [30, Theorem 3.1] and [31, Corollary 2.13]):
Theorem 5.1.
Let $V\subset\mathbb{C}^{n}$ be a complex algebraic set. Then the following statements are equivalent:
  • (1)
    $V$ is $\alpha$-Hölder smooth at infinity for all $\alpha\in(0,1)$;
  • (2)
    $V$ is strongly Hölder smooth at infinity;
  • (3)
    $V$ is log-Lipschitz smooth at infinity;
  • (4)
    $V$ is Lipschitz smooth at infinity;
  • (5)
    $V$ is the union of an affine linear subspace of $\mathbb{C}^{n}$ and a (possibly empty) finite set.
We also present a generalization of [3, Thereom II] for complex analytic sets in general. Indeed, we prove that a complex analytic set $V\subset\mathbb{C}^{n}$ is Lipschitz smooth at infinity if and only if $V$ is the union of an affine linear subspace of $\mathbb{C}^{n}$ and a (possibly empty) finite set (see Corollary 5.2). This gives a Lipschitz characterization of $\mathbb{C}^{d}$ as an affine linear subspace among all the pure-dimensional entire complex analytic sets.
The main ingredients in our proofs are the use of the new metric algebraic topology theory called moderately discontinuous homology introduced in [16] (see Subsection 2.4) and the use of a distance called here diameter distance. The diameter distance on a path-connected set $X$, which is defined as follows: given two points $x_{1},x_{2}\in X$, $d_{X,diam}(x_{1},x_{2})$ is the infimum of the diameters of the image of paths on $X$ connecting $x_{1}$ to $x_{2}$.
In general, one has that $|x_{1}-x_{2}|\leq d_{X,diam}(x_{1},x_{2})\leq d_{X,inn}(x_{1},x_{2})$, for all pair of points $x_{1},x_{2}\in X$, where $d_{X,inn}$ is the inner distance on $X$, which is defined as follows: given two points $x_{1},x_{2}\in X$, $d_{X,inn}(x_{1},x_{2})$ is the infimum of the lengths of paths on $X$ connecting $x_{1}$ to $x_{2}$.
An important class of sets where these distances are equivalent are the LNE sets. In order to know, we say that a path connected set $X\subset\mathbb{R}^{N}$ is Lipschitz normally embedded (LNE) if there exists a constant $C\geq 1$ such that $d_{X,inn}(x_{1},x_{2})\leq C\|x_{1}-x_{2}\|$, for all pair of points $x_{1},x_{2}\in X$.
However, these three distances are not equivalent in general, as it is shown in the next example:
Example 1.8.
Let $G$ be the graph of the function $f\colon\mathbb{R}\to\mathbb{R}$ given by
$f(x)=\left\{\begin{array}[]{ll}x\sin(\frac{1}{x}),&\mbox{ if }x\not=0,\\ 0,&\mbox{ if }x=0.\end{array}\right.$
Let $X=\{(tx,ty,t)\in\mathbb{R}^{3};t\geq 0$ and $(x,y)\in G\}$. For any point $(x,y)\in G\setminus\{(0,0)\}$, we have $d_{X,inn}((x,y,1),(0,0,1))=1+|(x,y)|\geq 2$. However, $d_{X,diam}((x,y,1),(0,0,1))\to 0$ as $(x,y)\to(0,0)$. Thus, $d_{X,diam}$ is not equivalent to $d_{X,inn}$.
In the case of the (real) cusp $C=\{(x,y)\in\mathbb{R}^{2};y^{3}=x^{2}\}$, which is a classical example of a set that is not LNE, we have that $d_{C,inn}$ and $d_{C,diam}$ are equivalent distances. Thus, we have the following natural problem:
Problem 2.
Let $X\subset\mathbb{R}^{n}$ be a path-connected set. Assume that $X$ is a definable set in an o-minimal structure on $\mathbb{R}$. Is $d_{X,diam}$ equivalent to $d_{X,inn}$?
In this article, we give a positive answer to Problem 2 (see Theorem 4.2).
Acknowledgements. The author would like to thank Ederson Braga for presenting the problem on log-Lipschitz regularity of complex analytic sets.
2. Preliminaries
All the subsets of $\mathbb{R}^{n}$ or $\mathbb{C}^{n}$ considered in the paper are supposed to be equipped with the Euclidean distance. When we consider other distance, it is clearly emphasized.
Notation:
  • $\|(x_{1},...,x_{n})\|=(x_{1}^{2}+...+x_{n}^{2})^{\frac{1}{2}}$;
  • $\mathbb{S}_{r}^{n-1}(p)=\{x\in\mathbb{R}^{n};\|x-p\|=r\}$, $\mathbb{S}_{r}^{n-1}=\mathbb{S}_{r}^{n-1}(0)$ and $\mathbb{S}^{n-1}=\mathbb{S}_{1}^{n-1}(0)$;
  • $B_{r}^{n}(p)=\{x\in\mathbb{R}^{n};\|x-p\|<r\}$;
  • For $X,Y\in\mathbb{R}^{n}$, $dist(X,Y):=\inf\{\|x-y\|;x\in X$ and $y\in Y\}$;
  • Let $f,g\colon(0,\varepsilon)\to[0,+\infty)$ be functions. We write $f\lesssim g$ if there is a constant $f(t)\leq Cg(t)$ for all $t\in(0,\varepsilon)$. We write $f\approx g$ if $f\lesssim g$ and $g\lesssim f$. We write $f\ll g$ if $\lim\limits_{t\to 0^{+}}\frac{f(t)}{g(t)}=0$.
2.1. Definitions of regularity
Definition 2.1.
Let $X\subset\mathbb{R}^{n}$ and $Y\subset\mathbb{R}^{m}$ be two sets and let $h\colon X\to Y$.
  • We say that $h$ is Lipschitz (resp. $\alpha$-Hölder) if there exists a positive constant $C$ such that $\|h(x)-h(y)\|\leq C\|x-y\|$ (resp. $\|h(x)-h(y)\|\leq C\|x-y\|^{\alpha}$) for all $x,y\in X$. In this case, we say that $C$ is a Lipschitz (resp. Hölder) constant of $h$.
  • We say that $h$ is bi-Lipschitz (resp. bi-$\alpha$-Hölder) if $h$ is a homeomorphism, it is Lipschitz (resp. $\alpha$-Hölder) and its inverse is also Lipschitz (resp. $\alpha$-Hölder). In this case, we also say that $X$ is bi-Lipschitz (resp. bi-$\alpha$-Hölder) homeomorphic to $Y$.
  • We say that $h$ is log-Lipschitz if there exists a positive constant $C$ such that $\|h(x)-h(y)\|\leq C\|x-y\|\big{|}\log\|x-y\|\big{|}$ for all $x,y\in X$ such that $20<\|x-y\|<1/$ 2.
  • We say that $h$ is bi-log-Lipschitz if $h$ is a homeomorphism, it is log-Lipschitz and its inverse is also log-Lipschitz. In this case, we also say that $X$ is bi-log-Lipschitz homeomorphic to $Y$.
Definition 2.2.
Let $X\subset\mathbb{R}^{n}$ be a set and $p\in X$.
  • We say that $X$ is $\alpha$-Hölder regular at $p$ if there exist $k\in\mathbb{N}$, $r>0$, an open neighbourhood $U\subset\mathbb{R}^{n}$ of $p$ and a bi-$\alpha$-Hölder homeomorphism $h\colon X\cap U\to B_{r}^{k}(0)$ such that $h(p)=0$.
  • We say that $X$ is strongly Hölder regular at $p$ if there exist $k\in\mathbb{N}$, $r>0$, an open neighbourhood $U\subset\mathbb{R}^{n}$ of $p$ and a homeomorphism $h\colon X\cap U\to B_{r}^{k}(0)$ that is a bi-$\alpha$-Hölder homeomorphism for all $\alpha\in(0,1)$ and such that $h(p)=0$.
  • We say that $X$ is Lipschitz (resp. log-Lipschitz) regular at $p$ if there exist $k\in\mathbb{N}$, $r>0$, an open neighbourhood $U\subset\mathbb{R}^{n}$ of $p$ and a bi-Lipschitz (resp. bi-log-Lipschitz) homeomorphism $h\colon X\cap U\to B_{r}^{k}(0)$ such that $h(p)=0$.
Definition 2.3.
Let $X\subset\mathbb{R}^{N}$ be a subset. We say that $X$ is Lipschitz normally embedded (LNE) if there exists a constant $C\geq 1$ such that $d_{X,inn}(x_{1},x_{2})\leq C\|x_{1}-x_{2}\|$, for all pair of points $x_{1},x_{2}\in X$. In this case, we also say that $X$ is $C$-LNE. We say that $X$ is LNE at $p\in\mathbb{R}^{N}$ if there exists an open neighborhood $U$ of $p$ such that $X\cap U$ is LNE. We say that $X$ is LNE at infinity if there exists a compact subset $K\subset\mathbb{R}^{N}$ such that $X\setminus K$ is LNE.
2.2. Tangent cones
Definition 2.4.
Let $X\subset\mathbb{R}^{n}$ be a subset and let $p\in\overline{X}$ (resp. $p=\infty$). We say that $v\in\mathbb{R}^{n}$ is a tangent vector of $X$ at $p$ ( resp. infinity) if there are a sequence $\{x_{j}\}_{j\in\mathbb{N}}\subset X$ and a sequence of positive real numbers $\{t_{j}\}_{j\in\mathbb{N}}$ such that $\lim\limits_{j\to\infty}t_{j}=0$ (resp. $\lim\limits_{j\to\infty}t_{j}=+\infty$) and $\lim\limits_{j\to\infty}\frac{1}{t_{j}}(x_{j}-p)=v$ (resp. $\lim\limits_{j\to\infty}\frac{x_{j}}{t_{j}}=v$). The set of all tangent vectors of $X$ at $p$ (resp. infinity) is denoted by $C(X,p)$ (resp. $C(X,\infty)$) and is called the tangent cone of $X$ at $p$ (resp. the tangent cone of $X$ at infinity).
Remark 2.5.
Let $X\subset\mathbb{C}^{n}$ be a complex analytic (resp. algebraic) set and $x_{0}\in X$. In this case, $C(X,x_{0})$ (resp. $C(X,\infty)$) is the zero set of a set of complex homogeneous polynomials, see this in [35, Theorem 4D] (resp. [22, Theorem 1.1] or [32, Theorem 3.1]). In particular, $C(X,x_{0})$ (resp. $C(X,\infty)$) is the union of complex lines passing through the origin $0\in\mathbb{C}^{n}$.
When $X\subset\mathbb{R}^{N}$ is a subanalytic (resp. semialgebraic) subset, we denote by $Link_{p}(X)$ (resp. $Link_{\infty}(X)$) to be the link of $X$ at $p$ (resp. infinity).
2.3. Pancake decomposition
Let $A\subset\mathbb{R}^{n}$ be a connected compact subanalytic set. By [19, Proposition 3], given $\epsilon>0$, there is a partition into a subanalytic and finite union $A=\bigcup_{i\in I}{B_{i}}$ such that every subset $B_{i}$ is $(1+\epsilon)$-LNE. For each $i$, let $X_{i}$ be the closure of $B_{i}$ in $A$. We consider $x,y\in A$ and for each $r$ and $k$ we define
$\widetilde{\Delta}_{r}(x,y):=\inf\left\{\sum_{i=0}^{r-1}{\|x_{i+1}-x_{i}\|};x_{0}=x,x_{r}=y,x_{i},x_{i+1}\in X_{\nu_{i}},0\leq i\leq r-1\right\}.$
$\Delta_{k}(x,y):=\inf\left\{\widetilde{\Delta}_{r}(x,y);r=1,\cdots,k\right\}.$
We also define $\inf\emptyset=+\infty$. It is clear that every $\Delta_{k}$ is subanalytic. Finally, we define
$d_{A,P}(x,y):=\inf\left\{\Delta_{k}(x,y);k\in\mathbb{N}\right\}.$
Proposition 2.6 (Lemma 4 and Theorem 1 in [19]).
The function $d_{A,P}\colon A\times A\rightarrow\mathbb{R}$ is subanalytic, defines a distance in $A$ and
$d_{A,P}(x,y)\leq d_{A,inn}(x,y)\leq(1+\epsilon)d_{A,P}(x,y),$
for all $x,y\in A$.
The distance $d_{A,P}$ is called a pancake distance on $A$.
2.4. Moderately discontinuous homology
Since Moderately discontinuous homology is a very recent theory, we briefly recall the main definitions of it. For more details about this theory, we refer the reader to [16]. The theory works for a large class of distances, but here we only consider the outer distance $d_{out}$, that one induced by the Euclidean space.
For $n\in\mathbb{N}$, let $\Delta_{n}\subset\mathbb{R}^{n+1}$ denote the standard $n$-simplex (i.e, $\Delta_{n}:=\{(p_{0},\ldots,p_{n})\in\mathbb{R}_{\geq 0}^{n+1}:\sum_{i=0}^{n}p_{i}=1\}$) with the orientation induced by the standard orientation in $\mathbb{R}^{n+1}$ of the convex hull of $\Delta_{n}\cup 0$. For $0\leq k\leq n$, denote by $i_{n}^{k}:\Delta_{n-1}\mapsto\Delta_{n},(p_{0},\ldots,p_{n-1})\mapsto(p_{0},\ldots,p_{k-1},0,p_{k},\ldots,p_{n-1})$. Set
$\hat{\Delta}_{n}:=\{(tx,t)\in\mathbb{R}^{n+1}\times\mathbb{R}:x\in\Delta_{n},t\in[0,1)\}$
and let $\hat{j}_{n}^{k}:\hat{\Delta}_{n-1}\to\hat{\Delta}_{n},(tx,t)\mapsto(ti_{n}^{k}(x),t)$. We identify $\hat{\Delta}_{n}$ with its germ at $(0,0)$.
Definition 2.7.
Given a definable germ $(X,x_{0})$. A linearly vertex approaching $n$-simplex (l.v.a. $n$-simplex) in $(X,x_{0})$ is a subanalytic continuous map germ : $\sigma:\hat{\Delta}_{n}\to(X,x_{0})$ such that there is $K\geq 1$ satisfying
$\frac{1}{K}t\leq\|\sigma(tx,t)-x_{0}\|\leq Kt,\forall(x,t)\in\hat{\Delta}_{n}.$
A linear vertex approaching $n$-chain in $(X,x_{0})$ is a finite formal sum $\sum_{i\in I}a_{i}\sigma_{i}$ where $a_{i}\in\mathbb{Z}$ and $\sigma_{i}$ is a l.v.a. $n$-simplex in $(X,x_{0})$. Denote by $MDC^{pre}_{n}(X,x_{0})$ the abelian group of $n$-chains. The boundary of $\sigma$ is a formal sum of $(n-1)$-simplices defined as follows:
$\partial\sigma:=\sum_{k=0}^{n}(-1)^{k}\sigma\circ j_{n}^{k}.$
Definition 2.8.
A homological subdivision of $\hat{\Delta}_{n}$ is a finite collection $\{\rho_{i}\}_{i\in I}$ of injective l.v.a. map germs $\rho_{i}:\hat{\Delta}_{n}\to\hat{\Delta}_{n}$ such that there is a subanalytic triangulation $\alpha:|K|\to\hat{\Delta}_{n}$ with the following properties
  • $\alpha$ is compatible with faces of $\hat{\Delta}_{n}$;
  • the collection $\{T_{i}\}$ of maximal triangles of $\alpha$ is also indexed by $I$;
  • for any $i\in I$, $\rho_{i}(\hat{\Delta}_{n})=T_{i}$ and the map $\alpha^{-1}\circ\rho_{i}$ takes faces of $\hat{\Delta}_{n}$ to faces of $K$.
The sign of $\rho_{i}$, denoted by $sign(\rho_{i})$, is defined to be $1$ if $\rho_{i}$ is orientation preserving, and $-1$ if it is of the opposite orientation.
Definition 2.9.
Let $b\in(0,+\infty]$. Two $n$-simplices $\sigma_{1}$ and $\sigma_{2}$ in $MDC_{n}^{pre}(X,x_{0})$ are called $b$-equivalent (write $\sigma_{1}\sim_{b}\sigma_{2}$) if
  • for $b<\infty$:
    $\lim_{t\to 0}\frac{\max\{d_{1}(\sigma_{1}(tx,t),\sigma_{2}(tx,t)),x\in\Delta_{n}\}}{t^{b}}=0.$
  • for $b=\infty$: $\sigma_{1}(tx,t)=\sigma_{2}(tx,t),\forall x\in\Delta_{n}$.
Definition 2.10.
Let $b\in(0,+\infty]$. Let $z=\sum_{i\in I}a_{i}\sigma_{i}$ and $z^{\prime}=\sum_{j\in J}b_{j}\tau_{j}$ be l.v.a. $n$-complex chains in $MDC_{n}^{pre}(X,x_{0})$. Write $I=\bigsqcup_{k\in K}I_{k}$ and $J=\bigsqcup_{k\in K}J_{k}$ where
(a) $i_{1},i_{2}\in I$ belong to the same $I_{k}$ iff $\sigma_{i_{1}}\sim_{b}\sigma_{i_{2}}$.
(b) $j_{1},j_{2}\in J$ belong to the same $J_{k}$ iff $\tau_{j_{1}}\sim_{b}\tau_{j_{2}}$.
(c) for $k\in K$, $i\in I_{k}$, $j\in J_{k}$, we have $\sigma_{i}\sim_{b}\tau_{j}$.
Then, $z$ is called $b$-equivalent to $z^{\prime}$, denoted $z\sim_{b}z^{\prime}$, if for any $k$,
$\sum_{i\in I_{k}}a_{i}=\sum_{j\in J_{k}}b_{j}.$
Definition 2.11.
Let $b\in(0,+\infty]$. Given two l.v.a. $n$-complex chains $z=\sum_{i\in I}a_{i}\sigma_{i}$ and $z^{\prime}=\sum_{j\in J}b_{j}\tau_{j}$ in $MDC_{n}^{pre}(X,x_{0})$.
  • We write $z\rightarrow_{b}z^{\prime}$ (immediate relation) if for each $i\in I$ there is a homological subdivision $\{\rho_{i,k}\}_{k\in K_{i}}$ such that
    $\sum_{i\in I}\sum_{k\in K_{i}}sgn(\rho_{i,k})a_{i}\sigma\circ\rho_{i,k}\sim_{b}\sum_{j\in J}b_{j}\tau_{j}.$
  • $z$ and $z^{\prime}$ are called homological subdivision equivalent, denoted by $z\sim_{S,b}z^{\prime}$, if there exist sequences of immediate sequences $z=z_{0}\rightarrow_{b}z_{1}\rightarrow_{b}\ldots\rightarrow_{b}z_{l}$ and $z^{\prime}=w_{0}\rightarrow_{b}w_{1}\rightarrow_{b}\ldots\rightarrow_{b}w_{m}$ such that $w_{m}\sim_{b}z_{l}$.
Definition 2.12.
Let $b\in(0,+\infty]$. The $b$-moderately discontinuous chain complex in $(X,x_{0})$ is the quotient group $MDC_{\bullet}^{b}(X,x_{0}):=MDC_{\bullet}^{pre}(X,x_{0})/\sim_{S,b}$. Its homology is called $b$-moderately discontinuous homology, denoted by $MDH_{\bullet}^{b}(X,x_{0})$.
Definition 2.13.
Let $X$ be a subanalytic germ at 00 in $\mathbb{R}^{n}$. Let $b\in(0,+\infty)$. The $b$-horn neighborhood of amplitude $\eta$ of $X$ is the subset
$\mathcal{H}_{b,\eta}(X):=\bigcup_{x\in X}\{z\in\mathbb{R}^{n};\|z-x\|<\eta\|x\|^{b}\}.$
The $\infty$-horn neighborhood $\mathcal{H}_{\infty,\eta}(X)$ is defined to be $X$.
We have the following two interesting results:
Theorem 2.14 (Theorem 13.5 in [16]).
Let $(X,x_{0})\subset(\mathbb{R}^{m},x_{0})$ be a closed subanalytic germ. Then there is a natural isomorphism
$MDH^{1}_{*}(X,x_{0})\to MDH^{1}_{*}(C(X,x_{0}),0)\to H_{*}(C(X,x_{0})\setminus\{0\}).$
Theorem 2.15.
Let $X$ be a subanalytic germ at 00 in $\mathbb{R}^{n}$. Let $b\in(0,+\infty)$ and $\eta>0$. There exists a $b^{\prime}$ satisfying $b<b^{\prime}$ such that the $b$-MD homology $MDH^{b}_{\bullet}(X,0)$ is isomorphic, for any $b^{\prime\prime}\in(b,b^{\prime})$, to the singular homology of the punctured $b^{\prime\prime}$-horn neighborhood $\mathcal{H}_{b^{\prime\prime},\eta}(X)\setminus\{0\}$.
3. Main result
Theorem 3.1.
Let $X\subset\mathbb{C}^{n}$ be a complex analytic set. Then the following statements are equivalent:
  • (1)
    $X$ is $\alpha$-Hölder regular at 00 for all $\alpha\in(0,1)$;
  • (2)
    $X$ is strongly Hölder regular at 00;
  • (3)
    $X$ is log-Lipschitz regular at 00;
  • (4)
    $X$ is Lipschitz regular at 00;
  • (5)
    $X$ is smooth at 00.
Proof.
It is clear that $(5)\Rightarrow(4)\Rightarrow(3)\Rightarrow(2)\Rightarrow(1)$. So, we only have to prove that $(1)\Rightarrow(5)$. In order to do that, we assume that $X$ is $\alpha$-Hölder regular at 00 for all $\alpha\in(0,1)$. Let $d$ be the dimension of $X$.
Let $\phi\colon X\to B$ be a bi-$\alpha$-Hölder homeomorphism, where $B$ is an open ball of $\mathbb{C}^{d}$. By McShane’s Theorem (see [24, Corollary 1]), there exist $\alpha$-Hölder maps $\widetilde{\phi}\colon\mathbb{C}^{n}\to\mathbb{C}^{d}$ and $\widetilde{\psi}\colon\mathbb{C}^{d}\to\mathbb{C}^{n}$ such that $\widetilde{\phi}|_{X}=\phi$ and $\widetilde{\psi}|_{Y}=\phi^{-1}$. Let us define $\varphi,\psi\colon\mathbb{C}^{n}\times\mathbb{C}^{d}\to\mathbb{C}^{d}\times\mathbb{C}^{n}$ as follows:
$\varphi(x,y)=(x-\widetilde{\psi}(y+\widetilde{\phi}(x)),y+\widetilde{\phi}(x))$
and
$\psi(z,w)=(z+\widetilde{\psi}(w),w-\widetilde{\phi}(z+\widetilde{\psi}(w))).$
It is easy to check that $\psi$ and $\varphi$ are $\alpha^{2}$-Hölder maps, $\psi=\varphi^{-1}$ and $\varphi(X\times\{0\})=\{0\}\times B$.
Thus, by doing the identification: $X\leftrightarrow X\times\{0\}$ one can suppose that for each $\alpha\in(0,1)$ there exists a bi-$\alpha$-Hölder homeomorphism $\varphi\colon(\mathbb{C}^{N},0)\to(\mathbb{C}^{N},0)$ such that $\varphi(X)=\{0\}\times B$, where $N=n+d$.
Let $f_{1},...,f_{k}\in\mathbb{C}[z_{1},...,z_{N}]$ be homogeneous polynomials such that $C(X,0)=V(f_{1},...,f_{k})=\{z\in\mathbb{C}^{N};f_{1}(z)=...=f_{k}(z)=0\}$. By changing $X$ by $X\times\mathbb{C}^{M}$, for some $M>0$, we may assume that $N\geq k+4$.
Let $X_{0}=C(X,0)\cap\mathbb{S}^{2N-1}$, and for $t>0$ let $X_{t}=(\frac{1}{t}X)\cap\mathbb{S}^{2N-1}$.
Claim 3.1.1.
$X_{0}$ is a connected set.
Proof of Claim 3.1.1.
Since $Link_{0}(X)$ is a connected set, for any $t>0$ smaller than the Milnor radius of $X$ at 00, $\varepsilon_{0}$, $X_{t}$ is a connected set.
We know that $\lim X_{t}=X_{0}$ with respect to the Hausdorff limit. It is well-known that the Hausdorff limit of a family of connected sets is a connected set as well. Since we could not find a reference for it, we present here a proof that $X_{0}$ is a connected set. Indeed, let $(A,B)$ be a separation of $X_{0}$, i.e., $A$ and $B$ are disjoint closed subsets of $X_{0}$ such that $X_{0}=A\cup B$. Since $X_{0}$ is a compact subset, we have that $A$ and $B$ are disjoint compact subsets, and thus we can find $\varepsilon>0$ such that $A_{\varepsilon}=\{x\in\mathbb{C}^{N};dist(x,A)<\varepsilon\}$ and $B_{\varepsilon}=\{x\in\mathbb{C}^{N};dist(x,B)<\varepsilon\}$ are disjoint open subsets. Thus, there exists $0<t_{0}\leq\varepsilon_{0}$ such that $X_{t}\subset U=A_{\varepsilon}\cup B_{\varepsilon}$ for all $0<t\leq t_{0}$. Then $(A_{\varepsilon}\cap X_{t},B_{\varepsilon}\cap X_{t})$ is a separation of $X_{t}$ for all $0<t\leq t_{0}$. Since $X_{t}$ is a connected set, $X_{t}\subset A_{\varepsilon}$ or $X_{t}\subset B_{\varepsilon}$ for all $0<t\leq t_{0}$. This shows that $A$ or $B$ is the empty set. Therefore $X_{0}$ is a connected set. ∎
Claim 3.1.2.
$H_{i}(X_{0})=0$ for all $21\leq i\leq 2d-$ 2.
Proof of Claim 3.1.2.
In order to prove this claim, we are going to use the moderately discontinuous homology. By Theorem 2.14, there is an isomorphism between $MDH^{1}_{*}(X,0)$ and the singular homology $H_{*}(X_{0})$. Then, we have to prove that $MDH^{1}_{i}(X,0)=0$ for all $21\leq i\leq 2d-$ 2.
By Theorem 2.15, for each $\eta>0$, there exists a $b^{\prime}$ satisfying $1<b^{\prime}$ such that the $1$-MD homology $MDH^{1}_{\bullet}(X,0)$ is isomorphic, for any $b^{\prime\prime}\in(b,b^{\prime})$, to the singular homology of the punctured $b^{\prime\prime}$-horn neighborhood $\mathcal{H}_{b^{\prime\prime},\eta}(X)\setminus\{0\}$. In fact, $b^{\prime}$ does not depend on $\eta$ (see [16, Remark 11.11]). Fix $i\in\{1,...,2d-2\}$ and take $\xi\in H_{i}(Y\setminus\{0\})$, where $Y=\mathcal{H}_{b^{\prime\prime},\eta}(X)$. Let $\varphi\colon(\mathbb{C}^{N},0)\to(\mathbb{C}^{N},0)$ be a bi-$\alpha$-Hölder homeomorphism such that $\varphi(X)=\{0\}\times B$, for some open ball $B\subset\mathbb{C}^{d}$ centered at 00. Let $C>0$ be a Hölder constant of $\varphi$ and $\varphi^{-1}$. Then $\varphi(|\xi|)\subset\mathcal{H}_{\alpha^{2}b^{\prime\prime},C\eta^{\alpha}}(\{0\}\times B)\setminus\{0\}$. Since $\mathcal{H}_{\alpha^{2}b^{\prime\prime},C\eta^{\alpha}}(\{0\}\times B)\setminus\{0\}$ and $B\setminus\{0\}$ has the same homotopy type and $H_{i}(B\setminus\{0\})=0$, there is a chain $\tilde{\tau}$ in $\mathcal{H}_{\alpha^{2}b^{\prime\prime},C\eta^{\alpha}}(\{0\}\times B)\setminus\{0\}$ such that $\partial\tilde{\tau}=\varphi_{*}(\xi)$. Therefore $\xi$ is homologous to zero in $\mathcal{H}_{\alpha^{4}b^{\prime\prime},C^{1+\alpha}\eta^{\alpha^{2}}}(X)\setminus\{0\}$.
By taking $\alpha$ such that $\alpha^{4}b^{\prime\prime}\in(1,b^{\prime})$, we obtain that $\xi$ is homologous to zero in $\mathcal{H}_{b^{\prime\prime},\eta}(X)\setminus\{0\}$. Therefore $H_{i}(\mathcal{H}_{b^{\prime\prime},\eta}(X)\setminus\{0\})=0$, and thus $H_{i}(X_{0})=0$. ∎
Therefore, $\widetilde{H}_{i}(X_{0})=0$ for all $20\leq i\leq 2d-$ 2. In particular, $H^{i}(X_{0})=0$ for all $21\leq i\leq 2d-$ 2.
Let $\pi\colon V^{*}\to\widetilde{V}$ denote the quotient map by the $\mathbb{C}^{*}$-action, where $V=C(X,0)$ and $V^{*}=V\setminus\{0\}$.
Thus, we have the following generalized version of Prill’s Theorem (see [26]).
Claim 3.1.3.
$V$ is a linear subspace.
Proof of Claim 3.1.3.
Remind that $V=V(f_{1},...,f_{k}):=\{z\in\mathbb{C}^{N};f_{1}(z)=...=f_{k}(z)=0\}$ and $N\geq k+4$.
Let $\mathbb{P}=\mathbb{C}P^{N-1}$ and $U=\mathbb{P}\setminus\widetilde{V}$. By the Alexander Duality,
$H^{s}(\mathbb{P},\widetilde{V})=H_{2(N-1)-s}(U).$
We are going to prove that $H_{m}(U)=0$ for all $m\geq N-1+k$. We proceed by induction on $k$. Indeed, if $k=1$, then this follows from Andreotti–Frankel theorem (see [2]), since $\mathbb{P}\setminus\mathbb{P}V(f_{1})$ is a smooth affine hypersurface of complex dimension $N-1$.
Assume that $H_{m}(\mathbb{P}\setminus\mathbb{P}V(g_{1},...,g_{k-1}))=0$ for all $m\geq N-1+k-1$, and any homogeneous polynomials $g_{1},...,g_{k-1}\in\mathbb{C}[z_{1},...,z_{N}]$. We have that $U=U^{\prime}\cup U_{k}$, where $U_{k}=\mathbb{P}\setminus\mathbb{P}V(f_{k})$ and $U^{\prime}=\mathbb{P}\setminus\mathbb{P}V(f_{1},...,f_{k-1})$. By hypothesis of induction, $H_{m}(\mathbb{P}\setminus\mathbb{P}V(f_{k}f_{1},...,f_{k}f_{k-1}))=0$ for all $m\geq N-1+k-1$. By the Mayer-Vietoris sequence, we obtain $H_{m}(\mathbb{P}\setminus\mathbb{P}V(f_{1},...,f_{k}))=0$ for all $m\geq N-1+k$.
Since $N\geq k+4$, then $H^{2}(\mathbb{P},\widetilde{V})=H^{3}(\mathbb{P},\widetilde{V})=0$, and thus the morphism $j^{2}\colon H^{2}(\mathbb{P})\to H^{2}(\widetilde{V})$, induced by the inclusion $j\colon\widetilde{V}\to\mathbb{P}$, is an isomorphism.
We know that the integral cohomology algebra $H^{\bullet}(\mathbb{P})$ is a truncated polynomial algebra $H^{\bullet}(\mathbb{P})=\mathbb{Z}[\alpha]/(\alpha^{N})$ generated by an element $\alpha$ of degree 2 (see [8, Chapter 5, Proposition 1.6]). Thus, $\alpha_{V}=j^{2}(\alpha)$ is a generator of $H^{2}(\widetilde{V})$.
The associated Gysin sequence in cohomology lead to the following
${...\rightarrow H^{2p+1}(X_{0})\rightarrow H^{2p}(\widetilde{V})}$${H^{2p+2}(\widetilde{V})\rightarrow H^{2p+2}(X_{0})\rightarrow...}$$\scriptstyle{\psi_{p}}$
where $\psi_{p}$ is the cup product with $\alpha_{V}$. Since $H^{i}(X_{0})=0$ for all $21\leq i\leq 2d-$ 2, we obtain $\psi_{p}$ is an isomorphism for all $p\in\{1,...,d-2\}$ and, in particular, $H^{2d-2}(\widetilde{V})=\mathbb{Z}$ and $\alpha_{V}^{d-1}$ is a generator of $H^{2d-2}(\widetilde{V})$. This shows also that $\widetilde{V}$ is an irreducible algebraic set.
Let $E_{m}$ be a generic hyperplane in $\mathbb{P}$ of dimension $m$. Then ${\rm deg}(\widetilde{V})=\#(\widetilde{V}\cap E_{N-d})$ and $[\widetilde{V}]={\rm deg}(\widetilde{V})[E_{d-1}]$ in $H^{2(d-1)}(\mathbb{P})$, where $[A]$ denotes the fundamental class of $A$. Hence,
$\langle j^{2(d-1)}(\alpha^{d-1}),[\widetilde{V}]\rangle=\langle\alpha_{V}^{d-1},[\widetilde{V}]\rangle=\langle\alpha_{V}^{d-1},{\rm deg}(\widetilde{V})[E_{d-1}]\rangle={\rm deg}(\widetilde{V}).$
Therefore $\alpha_{V}^{d-1}={\rm deg}(\widetilde{V})\cdot g$, where $g$ is a generator of $H^{2d-2}(\widetilde{V})$. Since $\alpha_{V}^{d-1}$ is also a generator of $H^{2d-2}(\widetilde{V})$, we obtain that ${\rm deg}(\widetilde{V})=1$, and therefore $V$ is a linear subspace.
Let us choose linear coordinates $(x,y)$ such that $C(X,0)=\{(x,y)\in\mathbb{C}^{N};y=0\}\cong\mathbb{C}^{d}$.
Let $P\colon\mathbb{C}^{N}\to\mathbb{C}^{d}$ be the mapping given by $P(x,y)=x$.
Claim 3.1.4.
There exist a constant $R>0$ and a polydisc $\Delta\subset\mathbb{C}^{N}$ centred at 00 such that $R\|P(x,y)\|\geq\|y\|$ for all $(x,y)\in X\cap\Delta$.
Proof of Claim 3.1.4.
Assume that Claim 3.1.4 is not true. Then there is a sequence $\{(x_{j},y_{j})\}_{j}\subset X$ such that $j\|x_{j}\|<\|y_{j}\|$ for all $j\in\mathbb{N}$ and $\lim\limits_{j\to+\infty}(x_{j},y_{j})=0$. In particular, $\lim\limits_{j\to+\infty}\frac{x_{j}}{\|(x_{j},y_{j})\|}=0$. By taking subsequence, if necessary, we may assume that $\lim\limits_{j\to+\infty}\frac{(x_{j},y_{j})}{\|(x_{j},y_{j})\|}=(0,v)$, with $\|v\|=1$. Then, $(0,v)\in P^{-1}(0)\cap C(X,0)$, which is a contradiction with $P^{-1}(0)\cap C(X,0)=\{0\}$.
Therefore Claim 3.1.4 holds true. ∎
Thus, $p=P|_{\Delta\cap X}\colon\Delta\cap X\to\Delta^{\prime}=\pi(\Delta\cap X)$ is a proper mapping. Then, there are a complex analytic set $\Sigma\subset\Delta^{\prime}$ with $\dim\Sigma<d$ and a positive integer number $k$ such that $p\colon\Delta\cap X\setminus p^{-1}(\Sigma)\to\Delta^{\prime}\setminus\Sigma$ is an analytic $k$-sheeted covering mapping.
Claim 3.1.5.
$k=1$.
Proof of Claim 3.1.5.
Let $v\in C(X,0)\setminus(\Sigma\cup C(\Sigma,0))$ be a unit vector. Then there exists $\varepsilon>0$ such that the arc $\beta\colon[0,\varepsilon)\to C(X,0)$ given by $\beta(t)=tv$ and satisfies $\beta(t)\not\in\Sigma\cup C(\Sigma,0)$ for all $t\in(0,\varepsilon)$. In fact, there exist positive real numbers $\eta$ and $\delta$ such that
$C_{\eta,\delta}=\{v\in C(X,\infty);\ \|v-tv_{0}\|<\eta t,\mbox{ for some }t\in(0,\delta)\}$
does not intersect the set $\Sigma$.
Assume that $k>1$. Then, there are two different arcs $\gamma_{1},\gamma_{2}\colon[0,\varepsilon)\to X$ such that $p\circ\gamma_{i}(t)=\beta(t)$ for all $t\in[0,\varepsilon)$, $i=1,2$.
Claim 3.1.6.
$b:=ord_{0}\|\gamma_{1}(t)-\gamma_{2}(t)\|>1$.
Proof of Claim 3.1.6.
We have to show that $\lim\limits_{t\to 0^{+}}\frac{\|\gamma_{1}(t)-\gamma_{2}(t)\|}{t}=0$.
In fact, let us write $\gamma_{i}(t)=(tv,y_{i}(t))$. By Claim 3.1.4, $\frac{\|y_{i}(t)\|}{t}\leq R$. So, the limit $\lim\limits_{t\to 0^{+}}\frac{y_{i}(t)}{t}$ exists, let us say, $\lim\limits_{t\to 0^{+}}\frac{y_{i}(t)}{t}=w_{i}$. Then, $\lim\limits_{t\to 0^{+}}\frac{\gamma_{i}(t)}{t}=(v,w_{i})\in C(X,0)$, for $i=1,2.$ Since $C(X,0)=\{(x,y)\in\mathbb{C}^{N};y=0\}$, we obtain that $w_{1}=w_{2}=0$, and thus $\lim\limits_{t\to 0^{+}}\frac{\gamma_{1}(t)}{t}=\lim\limits_{t\to 0^{+}}\frac{\gamma_{2}(t)}{t}$, which finishes the proof of Claim 3.1.6. ∎
Let $d_{X,P}$ be a pancake distance on $X$ which is equivalent to $d_{X,inn}$ (see Proposition 2.6). Since $d_{X,P}$ is a subanalytic distance, $ord_{0}d_{X,P}(\gamma_{1}(t),\gamma_{2}(t))$ is a well-defined positive rational number, and thus we define
$ord_{0}d_{X,inn}(\gamma_{1}(t),\gamma_{2}(t)):=ord_{0}d_{X,P}(\gamma_{1}(t),\gamma_{2}(t)).$
Claim 3.1.7.
$ord_{0}d_{X,inn}(\gamma_{1}(t),\gamma_{2}(t))=1$.
Proof of Claim 3.1.7.
Let $\beta_{t}\colon[0,1]\to X$ be an arc connecting $\gamma_{1}(t)$ and $\gamma_{2}(t)$. Thus, $p\circ\beta_{t}$ is a loop in $\mathbb{C}^{d}$ based at $tv$. Since $p\colon\Delta\cap X\setminus p^{-1}(\Sigma)\to\Delta^{\prime}\setminus\Sigma$ is a covering mapping and $C_{\eta,\delta}$ is simply connected subset of $\Delta^{\prime}\setminus\Sigma$, the image of $p\circ\beta_{t}$ can not be contained in $C_{\eta,\delta}$, and thus $\ell(p\circ\beta_{t})\geq 2\eta t$. By using the facts that $p$ is 1-Lipschitz and that $1$-Lipschitz mappings do not increase the length of arcs, we obtain
$\ell(\beta_{t})\geq\ell(p\circ\beta_{t})\geq 2\eta t.$
By taking the infimum over all $\beta_{t}$, we obtain $d_{X,inn}(\gamma_{1}(t),\gamma_{2}(t))\geq 2\eta t$. Therefore $ord_{0}d_{X,inn}(\gamma_{1}(t),\gamma_{2}(t))=1$. ∎
Let $\alpha\in(0,1)$ be a number such that $\alpha^{2}>\frac{1}{b}$.
By hypothesis, there are a neighbourhood of 00, $U\subset\mathbb{C}^{N}$, $r>0$ and a bi-$\alpha$-Hölder homeomorphism $\psi\colon B_{r}^{2d}(0)\to X\cap U$. Thus, there is $M>0$ such that
$\frac{1}{M}\|x-y\|^{\frac{1}{\alpha}}\leq\|\psi(x)-\psi(y)\|\leq M\|x-y\|^{\alpha},\quad\forall x,y\in B_{r}^{2d}(0).$
This implies that $d_{X,diam}(u,v)\leq M^{2}\|u-v\|^{\alpha^{2}}$ for all small enough $u,v\in X$.
It follows from [19, Proposition 3] (see also Subsection 2.3) that we can write $X=\bigcup\limits_{i=1}^{m}X_{i}$ such that each $X_{i}$ is an LNE closed set. Let $C$ be a positive number such that each $X_{i}$ is $C$-LNE for all $i\in\{1,...,m\}$.
Claim 3.1.8.
$d_{X,inn}(x,y)\leq(mC)\cdot d_{X,diam}(x,y)$ for all $x,y\in X$.
Proof of Claim 3.1.8.
Fix $x,y\in X$ and let $\alpha\colon[0,1]\to X$ be an arc connecting $x$ and $y$.
We define $s_{0}=0$. Let $i_{1}$ such that $\gamma_{1}=\alpha(0)\in X_{i_{1}}$. Let $s_{1}=\sup\{s\in[0,1];\alpha(s)\in X_{i_{1}}\}$. Since $X_{i_{1}}$ is a closed set, $\alpha(s_{1})\in X_{i_{1}}$. Assume that $s_{j}$ and $i_{j}$ are defined. If $s_{j}=1$, we stop the process. If $s_{j}<1$, let $i_{j+1}\not\in\{i_{1},...,i_{j}\}$ such that $\alpha(s_{j})\in X_{i_{j+1}}$. We define $s_{j+1}=\sup\{s\in[s_{j},1];\alpha(s)\in X_{i_{j+1}}\}$. Since $X_{i_{j+1}}$ is a closed set, $\alpha(s_{j+1})\in X_{i_{j+1}}$. Since we only have a finite number of $X_{i}$’s, there is a $j_{0}$ such that $s_{j_{0}}=1$. For each $j\in\{0,...,j_{0}-1\}$, we consider $\beta_{j}\colon[0,1]\to X_{i_{j+1}}$ be a minimizing geodesic connecting $\alpha(s_{j})$ and $\alpha(s_{j+1})$. Let $\beta\colon[0,1]\to X$ be the broken geodesic $\beta_{0}\cdot...\cdot\beta_{j_{0}-1}$.
Since $diam(Im(\alpha))\geq diam(Im(\alpha)\cap X_{i_{j+1}})\geq\|\alpha(s_{j+1})-\alpha(s_{j})\|$ for all $j\in\{0,...,j_{0}-1\}$, we obtain
$m\cdot diam(Im(\alpha))\geq j_{0}\cdot diam(Im(\alpha))\geq\frac{1}{C}\ell(\beta)\geq\frac{1}{C}d_{X,inn}(x,y).$
By taking the infimum over all $\alpha$, we obtain
$d_{X,inn}(x,y)\leq(mC)\cdot d_{X,diam}(x,y).$
Then $d_{X,inn}(\gamma_{1}(t),\gamma_{2}(t))\leq mCM^{2}\|\gamma_{1}(t)-\gamma_{2}(t)\|^{\alpha^{2}}$ for all small enough $t>0$, and thus $1\geq\alpha^{2}b$, which is a contradiction with the chosen $\alpha$.
Therefore $k=1$. ∎
Since $k=1$, we have that $p\colon\Delta\cap X\setminus p^{-1}(\sigma)\to\Delta^{\prime}\setminus\sigma$ is an analytic diffeomorphism. Let $\Phi\colon\Delta^{\prime}\setminus\Sigma\to X\setminus p^{-1}(\Sigma)$ be its inverse. Thus, there is an analytic mapping $\phi\colon\Delta^{\prime}\setminus\Sigma\to\mathbb{C}^{n-d}$ such that $\Phi(x)=(x,\phi(x))$ for all $x\in\Delta^{\prime}\setminus\Sigma$. By Claim 3.1.4, there exists a constant $R>0$ such that $R\|p(x,y)\|\geq\|y\|$ for all $(x,y)\in X\cap\Delta$, then $\phi$ is a bounded mapping on $\Delta^{\prime}\setminus\Sigma$. Since $\Sigma$ is a removable set, $\phi$ can be extended to a complex analytic mapping $\tilde{\phi}\colon\Delta^{\prime}\to\mathbb{C}^{n-d}$. Therefore, $X\cap\Delta$ is the graph of the analytic mapping $\tilde{\phi}$ and, in particular, $X$ is smooth at 00. ∎
The next result says that Theorem 3.1 is sharp in the sense that one cannot change the item $(1)$ in Theorem 3.1 by the following weaker statement: $X$ is $\alpha$-Hölder regular at 00, for some $\alpha\in(0,1)$ that only depends on the dimension of $X$.
Proposition 3.2.
For any $\alpha\in(0,1)$ and any positive integer $d$, there is a complex algebraic hypersurface $X\subset\mathbb{C}^{d+1}$ that is $\alpha$-Hölder regular at 00, but is not smooth at 00.
Proof.
For $\alpha\in(0,1)$, let $n>1$ be a integer such that $\alpha<\frac{n}{n+1}$.
Let $X=\{(x,y)\in\mathbb{C}^{2};x^{n}=y^{n+1}\}$ and $L=X\cap\mathbb{S}^{3}$. Since $X$ is an irreducible and weighted-homogeneous curve, then there is a diffeomorphism $f\colon\mathbb{S}^{1}\rightarrow L$. In particular, $f$ is a bi-Lipschitz homeomorphism, i.e., there is $\lambda\geq 1$ such that
$\frac{1}{\lambda}\|v_{1}-v_{2}\|\leq\|f(v_{1})-f(v_{2})\|\leq\lambda\|v_{1}-v_{2}\|\ \forall\ v_{1},v_{2}\in\mathbb{S}^{1}.$
We write $f(v)=(f_{1}(v),f_{2}(v))$ and define $F\colon\mathbb{C}\rightarrow X$ by
$F(tv)=(tf_{1}(v),t^{\frac{n}{n+1}}f_{2}(v))\quad\mbox{ for all }v\in\mathbb{S}^{1}\mbox{ and }t\geq 0.$
We claim that $F$ is a $b$-Hölder mapping around 00, where $b=\frac{n}{n+1}$. Indeed, given $u_{1}=tv_{1},u_{2}=sv_{2}\in\mathbb{C}$ with $v_{1},v_{2}\in\mathbb{S}^{1}$ and $t\leq s\leq 1$, we have
$\|F(u_{1})-F(u_{2})\|$$\leq$ $\|F(tv_{1})-F(tv_{2})\|+\|F(tv_{2})-F(sv_{2})\|$
$\leq$ $t\|f_{1}(v_{1})-f_{1}(v_{2})\|+t^{b}\|f_{2}(v_{1})-f_{2}(v_{2})\|$
$+|t-s|\|f_{1}(v_{2})\|+|t^{b}-s^{b}|\|f_{2}(v_{2})\|$
$\leq$ $\lambda t\|v_{1}-v_{2}\|+\lambda t^{b}\|v_{1}-v_{2}\|+|t-s|\|v_{2}\|+|t^{b}-s^{b}|\|v_{2}\|$
$\leq$ $(\lambda+1)\|tv_{1}-sv_{2}\|+2^{1-b}\lambda\|tv_{1}-tv_{2}\|^{b}+|t-s|^{b}\|v_{2}\|$
$\leq$ $(2\lambda+1)\|u_{1}-u_{2}\|+(2^{1-b}\lambda+1)\|tv_{1}-sv_{2}\|^{b}$
$\leq$ $(6\lambda+3)\|u_{1}-u_{2}\|^{b}.$
If we denote by $g\colon L\rightarrow\mathbb{S}^{1}$ the inverse mapping of $f$, we have that $G\colon X\rightarrow\mathbb{C}$ defined by $G(tx,t^{b}y)=tg(x,y)$ and $t\geq 0$ is the inverse of $F$ and it is a Lipschitz mapping around 00. Indeed, given $w_{1}=(tx_{1},t^{b}y_{2}),w_{2}=(sx_{2},s^{b}y_{2})\in X$ with $(x_{1},y_{1}),(x_{2},y_{2})\in L$ and $0\leq t\leq s\leq 1$, we have
$\|G(w_{1})-G(w_{2})\|$$\leq$ $t\|g(x_{1},y_{1})-g(x_{2},y_{2})\|+|t-s|\|g(x_{2},y_{2})\|$
$\leq$ $\lambda t\|(x_{1},y_{1})-(x_{2},y_{2})\|+|t-s|\|(x_{2},y_{2})\|$
$\leq$ $(\lambda+1)\|t(x_{1},y_{1})-s(x_{2},y_{2})\|$
$\leq$ $(\lambda+1)(\|tx_{1}-sx_{2}\|+\|ty_{1}-sy_{2}\|)$
$\leq$ $(\lambda+1)(1+\sqrt{2})\|w_{1}-w_{2}\|.$
Since $\alpha<b$, this implies that $F$ is a bi-$\alpha$-Hölder homeomorphism around 00, and thus $X$ is $\alpha$-Hölder regular at 00.
Moreover, for any positive integer $d$, $X\times\mathbb{C}^{d-1}\subset\mathbb{C}^{d+1}$ is $\alpha$-Hölder regular at 00, but is not smooth at 00.
4. The diameter and the inner distances are equivalent
Here, we prove a result that is slightly more general than Claim 3.1.8, since here we do not assume that the set is compact.
Proposition 4.1.
Let $X$ be a path-connected set. Let $X=\bigcup_{i\in I}{B_{i}}$ be a finite decomposition such that each $B_{i}$ is $C$-LNE. Then
$d_{X,diam}\leq(\#I\cdot C)d_{X,inn}.$
Proof.
Fix $\varepsilon>0$ and let $m=\#I$. For each $i$, let $X_{i}$ be the closure of $B_{i}$ in $X$.
Fix $x,y\in X$ and let $\alpha\colon[0,1]\to X$ be an arc connecting $x$ and $y$.
We define $s_{0}=0$. Let $i_{1}$ such that $\gamma_{1}=\alpha(0)\in X_{i_{1}}$. Let $s_{1}=\sup\{s\in[0,1];\alpha(s)\in X_{i_{1}}\}$. Since $X_{i_{1}}$ is a closed set in $X$, $\alpha(s_{1})\in X_{i_{1}}$. Assume that $s_{j}$ and $i_{j}$ are defined. If $s_{j}=1$, we stop the process. If $s_{j}<1$, let $i_{j+1}\not\in\{i_{1},...,i_{j}\}$ such that $\alpha(s_{j})\in X_{i_{j+1}}$. We define $s_{j+1}=\sup\{s\in[s_{j},1];\alpha(s)\in X_{i_{j+1}}\}$. Since $X_{i_{j+1}}$ is a closed set, $\alpha(s_{j+1})\in X_{i_{j+1}}$. Since we only have a finite number of $X_{i}$’s, there is a $j_{0}$ such that $s_{j_{0}}=1$. For each $j\in\{0,...,j_{0}-1\}$, we consider $\beta_{j}\colon[0,1]\to X_{i_{j+1}}$ be an arc connecting $\alpha(s_{j})$ and $\alpha(s_{j+1})$ such that $length(\beta_{j})\leq(1+\varepsilon)d_{X,inn}(\alpha(s_{j}),\alpha(s_{j+1}))$. Let $\beta\colon[0,1]\to X$ be the arc $\beta_{0}\cdot...\cdot\beta_{j_{0}-1}$.
Since $diam(Im(\alpha))\geq diam(Im(\alpha)\cap X_{i_{j+1}})\geq\|\alpha(s_{j+1})-\alpha(s_{j})\|$ for all $j\in\{0,...,j_{0}-1\}$, we obtain
$m\cdot diam(Im(\alpha))\geq j_{0}\cdot diam(Im(\alpha))\geq\frac{1}{C(1+\varepsilon)}\ell(\beta)\geq\frac{1}{C(1+\varepsilon)}d_{X,inn}(x,y).$
By taking the infimum over all $\alpha$, we obtain
$d_{X,inn}(x,y)\leq(mC(1+\varepsilon))\cdot d_{X,diam}(x,y)$
and setting $\varepsilon\to 0^{+}$, we have
$d_{X,inn}(x,y)\leq(mC)\cdot d_{X,diam}(x,y).$
Since any o-minimal set can be partitioned into a finite union of definable LNE sets (e.g., see [20, Theorem 1.3]), we obtain the following result, which gives a positive answer to Problem 2.
Theorem 4.2.
Let $X$ be a connected definable set in an o-minimal structure on $\mathbb{R}$. Then $d_{X,diam}$ is equivalent to $d_{X,inn}$.
5. A log-Lipschitz version of a result of Ahern and Rudin
Theorem 5.1.
Let $V\subset\mathbb{C}^{n}$ be a complex algebraic set. Then the following statements are equivalent:
  • (1)
    $V$ is $\alpha$-Hölder smooth at infinity for all $\alpha\in(0,1)$;
  • (2)
    $V$ is strongly Hölder smooth at infinity;
  • (3)
    $V$ is log-Lipschitz smooth at infinity;
  • (4)
    $V$ is Lipschitz smooth at infinity;
  • (5)
    $V$ is the union of an affine linear subspace of $\mathbb{C}^{n}$ and a (possibly empty) finite set.
Proof.
It is clear that $(5)\Rightarrow(4)\Rightarrow(3)\Rightarrow(2)\Rightarrow(1)$. So, we only have to prove that $(1)\Rightarrow(5)$.
Assume that $V$ is $\alpha$-Hölder smooth at infinity for all $\alpha\in(0,1)$. In this case, $V$ is the union of a pure dimensional algebraic set and a (possibly empty) finite set. Thus, we may assume that $V$ is a pure dimensional set.
Let $A=\iota(V\setminus\{0\})\cup\{0\}$, where $\iota\colon\mathbb{C}^{n}\setminus\{0\}\to\mathbb{C}^{n}\setminus\{0\}$ is the mapping given by $\iota(x)=\frac{x}{\|x\|^{2}}$. It is easy to check that $C(V,\infty)=C(A,0)$. Then $C(A,0)$ is a pure $d$-dimensional complex homogeneous algebraic set, where $d=\dim_{\mathbb{C}}V$. By Claims 3.1.1 and 3.1.2, $\widetilde{H}_{i}(A_{0})=0$ for all $i\in\{0,...,2d-2\}$, where $A_{0}=C(A,0)\cap\mathbb{S}^{2n-1}$.
By Claim 3.1.3, $C(A,0)=C(V,\infty)$ must be a complex linear subspace.
By following the proof of Theorem 3.1, if $V$ is not an affine linear subspace, then for a generic $v\in C(V,\infty)$ there are two different arcs $\gamma_{1},\gamma_{2}\colon(\varepsilon,\infty)\to V$ such that $p(\gamma_{j}(t))=tv$ for all $t\in(\varepsilon,\infty)$ and $j=1,2$, where $p$ is the restriction to $V$ of the orthogonal projection onto $C(V,\infty)$. Moreover,
$ord_{\infty}d_{V,inn}(\gamma_{1},\gamma_{2}):=\lim\limits_{t\to+\infty}\frac{\log d_{V,inn}(\gamma_{1}(t),\gamma_{2}(t))}{\log t}=1$
and $\lim\limits_{t\to+\infty}\frac{\gamma_{1}(t)-\gamma_{2}(t)}{t}=0$, and thus we have that
$b=ord_{\infty}\|\gamma_{1}-\gamma_{2}\|:=\lim\limits_{t\to+\infty}\frac{\log\|\gamma_{1}(t)-\gamma_{2}(t)\|}{\log t}<1.$
For each $j\in\{1,2\}$, let $\tilde{\gamma}_{j}\colon(0,\frac{1}{\varepsilon})\to A$ be the arc given by $\tilde{\gamma}_{j}(s)=\iota\circ\gamma\left(\frac{1}{s}\right)$. Then, $ord_{0}d_{A,inn}(\tilde{\gamma}_{1},\tilde{\gamma}_{2})=1$ and $\tilde{b}=ord_{0}\|\tilde{\gamma}_{1}-\tilde{\gamma}_{2}\|=2-b>1$ (see the proof of Corollary 5.7 in [12]).
Again, by following the proof of Theorem 3.1, for $\alpha\in(\tilde{b}^{-\frac{1}{2}},1)$, there is a constant $M_{\alpha}>0$ such that $d_{A,diam}(u,v)\leq M_{\alpha}\|u-v\|^{\alpha^{2}}$ for all small enough $u,v\in A$. Since $A$ is a path-connected semialgebraic set, by Theorem 4.2, there is a constant $C$ such that $d_{A,inn}\leq Cd_{A,diam}$. In particular, $d_{A,inn}(\tilde{\gamma}_{1}(s),\tilde{\gamma}_{2}(s))\leq CM_{\alpha}\|\tilde{\gamma}_{1}(t)-\tilde{\gamma}_{2}(t)\|^{\alpha^{2}}$ for all small enough $t>0$, and thus $11\geq\alpha^{2}\tilde{b}>$ 1, which is a contradiction. Therefore, $V$ must be an affine linear subspace. ∎
As a direct consequence, we obtain the next result, which generalizes [3, Thereom II].
Corollary 5.2.
A complex analytic set $V\subset\mathbb{C}^{n}$ is Lipschitz smooth at infinity if and only if $V$ is the union of an affine linear subspace of $\mathbb{C}^{n}$ and a (possibly empty) finite set.
Proof.
Assume that $V$ is Lipschitz smooth at infinity. From [33, Theorem 5.1] and [32, Theorem 3.1], we obtain that $V$ is a complex algebraic set. By Theorem 5.1, $V$ is the union of an affine linear subspace of $\mathbb{C}^{n}$ and a (possibly empty) finite set.
The reciprocal is obvious. ∎
Ahern and Rudin in [3] defined also the notion of a mapping to be $C^{1}$-smooth at infinity.
Definition 5.3.
A map $F\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ is $C^{1}$-smooth at infinity if $\iota\circ F\circ\iota$ extends to a $C^{1}$ mapping in a neighbourhood of the origin of $\mathbb{R}^{n}$.
Thus, we define also the following:
Definition 5.4.
A map $F\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ is Lipschitz smooth at infinity if $\iota\circ F\circ\iota$ extends to a Lipschitz mapping in a neighbourhood of the origin of $\mathbb{R}^{n}$.
It is also clear that if $F\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ is $C^{1}$-smooth at infinity then it is Lipschitz smooth at infinity. Thus, we obtain the following version of [3, Thereom I]:
Corollary 5.5.
A bi-holomorphism $F\colon\mathbb{C}^{n}\to\mathbb{C}^{n}$ is Lipschitz smooth at infinity if and only if $F$ is an affine mapping.
Proof.
It is clear that any affine mapping that is a bi-holomorphism must be a bi-Lipschitz homeomorphism. Then, by [33, Corollary 5.2], $\iota\circ F\circ\iota$ extends to a bi-Lipschitz homeomorphism $\tilde{F}\colon\mathbb{C}^{n}\to\mathbb{C}^{n}$. In particular, $F\colon\mathbb{C}^{n}\to\mathbb{C}^{n}$ is Lipschitz smooth at infinity.
Reciprocally, assume that $F\colon\mathbb{C}^{n}\to\mathbb{C}^{n}$ is Lipschitz smooth at infinity. By [33, Theorem 5.1], $F\colon\mathbb{C}^{n}\to\mathbb{C}^{n}$ is Lipschitz at infinity. By Liouville’s Theorem for holomorphic mappings, $F$ is an affine mapping. ∎